Solve each equation for x.
x^2 +4x+4=7
(Hint: Factor the left-hand side.)

Answer by checkley75(3666)

X^2+4X+4=7
X^2+4X+4-7=0
X^2+4X-3=0
USING THE QUADRATIC EQUATTION WE GET:
X=(-4+-SQRT[4^2-4*1*-3])/2*1
X=(-4+-SQRT[16+12])/2
X=(-4+-SQRT28)/2
X=(-4+-5.29)/2
X=(-4+5.29)/2
X=1.29/2
X=.645 ANSWER
X=(-4-5.29)/2
X=-9.29/2
X=-4.645 ANSWER.